# From Memories to Maps: Mechanisms of In-Context Reinforcement Learning in Transformers

### v2

@@ 4 Results
Gridworld A key signature of cross-context structure learning is the alignment of internal representations across environments with shared topology. Specifically, even when sensory observations differ across environments, states occupying the same grid location should be encoded more similarly than states from different locations (Fig. 4A). To test this, we compute pairwise Pearson correlation coefficients between model representations of the same graph node across different test environments (Fig. 4B). We separate correlation scores by whether the compared states occupy the same or different graph nodes. With limited context, representations appear collapsed—showing high correlation across all states regardless of node identity. As context length increases, this gap widens: states from the same node become more aligned than those from different nodes (Fig. 4B; solid vs. dashed).
@@
Tree Mazes Do similar alignment strategies emerge in models trained on tree mazes? As in gridworld, correlation scores across environments (Fig. 4D) reveal initially collapsed representations that differentiate by graph node with increasing context. We further stratify these correlations by node position within the tree (Fig. 4F). Cross-context alignment is strongest for states near the root or leaves of the tree, and weakest at intermediate depths. This pattern echoes the in-context results: the model favors encoding coarse structural motifs over precise positional detail. The PCA visualization (Fig. 4E) further illustrates this alignment across environments. Importantly, neither in-context nor cross-context representation alignment was explicitly trained—these strategies emerge opportunistically as a byproduct of meta-learning.
@@
Figure 5: Memory retrieval at decision time shows limited expansion from the query state and the goal state. A. Two example tree maze environments. The query state is marked with a red dot and the goal with a red star. Context length is 800. Edge color indicates gradient attribution strength for each transition, normalized within each environment. B. Average attribution strength as a function of distance from the query state, computed across 50 environments. Shaded regions show 95% confidence intervals. C. As in (B), but plotted as a function of distance from the goal state.
4.4 The RL strategy that emerges from meta-learning does not show signatures of value-based learning or model-based planning.
Thus far, we have shown that meta-learned agents replicate the rapid learning dynamics observed in animal behavior. A key component of this success is the emergence of structured representations from contextual input. We now turn to characterizing the mechanisms of the underlying RL strategy. We begin by testing whether the model exhibits hallmarks of standard model-free reinforcement learning. Specifically, we assess whether the model encodes value estimates $V^{*}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}\mid S_{0}=s\right]$ for its current state $s$ under an optimal policy. We train linear decoders on query token representations to predict $V(s)$ and evaluate generalization on held-out environments (Appendix J). In gridworld, $V(s)$ can be decoded with high accuracy (Appendix K). However, decoding is more accurate when $V(s)$ is defined in terms of Euclidean distance to the goal, rather than over the true 4-dimensional action space (Appendix K). This suggests that the model encodes spatial structure rather than true value gradients—its apparent $V(s)$ reflects geometric regularities, not action-contingent reward prediction. In tree mazes, decoded value estimates are localized: $V(s)$ is only reliable within 2–3 steps of the reward (Appendix K). This narrow value gradient is insufficient to guide behavior over the full task horizon.
Next, we test whether the model exhibits hallmarks of standard model-based RL. Such algorithms typically rely on explicit path planning between the query and goal states. Planning need not follow a strictly forward trajectory, and transformers in particular can implement diverse state-tracking strategies to support long-range planning [Li et al., 2025]. Critically, however, all such strategies depend on retrieving intermediate states along the path from query to goal during decision time. To evaluate this, we assess which context-memory tokens influence the model’s decision at a given query state. Using integrated gradients, we measure attribution strength for memory tokens along the query–goal path (Fig. 5; Appendix L). In both gridworld and tree mazes, tokens near the query and goal states show high attribution, while intermediate tokens receive little to no weight. This pattern is inconsistent with path planning, which requires attending to transitions along the full route at decision time (Appendix L).
As a further test, we perform attention ablations to assess the necessity of individual memory tokens for decision-making. To test whether information about a particular state $s$ is necessary, we mask attention from the query token to all context-memory tokens involving $s$. We repeat this procedure across model layers and measure the resulting change in cross-entropy loss (Appendix L). In both task suites, we find that the query token can make correct decisions without attending to intermediate path states. This again contradicts planning-based strategies, which require integration over intermediate transitions (Appendix L). Our conclusion contrasts with prior work that analyzed the behavior of meta-learned RL agents and suggested that they implement a form of model-based planning [Wang et al., 2016, Ritter et al., 2018]. It is possible that meta-learning can discover more typical model-based strategies in other settings, and these discrepancies demonstrate that the mechanistic perspective taken here can provide insights into understanding meta-learned algorithms.
@@
In summary, we find that the agent does not use value gradients or path planning to make decisions. Our analyses also reveal an additional neural prediction: memory retrieval at decision time should be biased toward experiences near the agent’s current location and its goal. Such replay patterns have been observed in the hippocampus during spatial decision-making tasks [Jackson et al., 2006, Pfeiffer and Foster, 2013, Mattar and Daw, 2018]. Overall, we suggest that the in-context strategy learned by the model lies outside the standard taxonomy of model-free and model-based reinforcement learning.
@@
Figure 6: Gridworld tasks are solved by aligning internal representations to Euclidean space. A. Overview of model analysis methods used in gridworld: linear decoding, attention ablation, and attention profiling. B. Total decoding error (X + Y) across $60$ test environments, plotted by true XY position of the query state. Errors shown for query stream representations from layer 1 (left) and layer 2 (right). Context length = 250. C. Decoded vs. true angle from query state to the goal across $60$ test environments. Results shown for representations from layer 1 (left) and layer 2 (right). D. Change in cross-entropy loss after ablating context tokens along the query–goal path, plotted by token position. Line color indicates layer of intervention. Mean across 50 environments; shaded regions show 95% confidence intervals. E. Average attention score between context-memory tokens as a function of spatial distance (query–key pairs), for layers 1 and 2. Mean over 50 environments; shaded regions show 95% confidence intervals.
@@
4.5 Transformers learn diverse in-context RL strategies where intermediate computations are stored in context-memory tokens.
@@
    Use in-context experience to align representations to Euclidean space, likely by locally stitching neighboring transitions.
@@
We arrived at this hypothesis by first identifying the task-relevant variables that are linearly decodable from the query token at each layer (Fig. 6A). As shown in Section 4.3, the model aligns states across environments based on shared grid structure. To test for spatial encoding, we train a linear decoder to predict the latent XY position of the query state from its embedding (Appendix J). Decoder accuracy is evaluated on held-out environments to test generalization across tasks. Decoding accuracy improves across layers, with spatial position becoming reliably recoverable by layer 2 (error $<1$; Fig. 6B). Estimation error varies spatially, with corner states decoded more accurately than interior locations (Fig. 6B). Building on this spatial structure, we next decode the angle from the query state to the goal. By layer 2, the angle to the goal is decoded with high accuracy (Fig. 6C). Strikingly, both position and goal angle can also be decoded from the context-memory tokens. By layer 2, individual context tokens—each representing a transition $(s,a,s^{\prime},r)$—encode not only the states involved, but also their positions and goal-relative angles.
To localize where angle-to-goal information is computed, we assess which context-memory tokens are necessary for correct decisions. We do this by masking attention from the query token to individual context tokens, following the ablation procedure described in Section 4.4. We find that layers 1 and 3 are robust to these ablations, but performance degrades in layer 2 when attention to tokens near the query or goal is removed (Fig. 6D). This suggests that layer 2 computes the relative angle between query and goal states, likely by comparing their internal XY coordinates. To understand how this spatial information becomes available for computation, we examine how state representations evolve through context-to-context attention. We analyze the spatial locality of attention between context-memory tokens to test whether transitions are integrated locally or globally. Specifically, we plot attention strength as a function of spatial distance between token pairs (Fig. 6E). We restrict this analysis to the first two layers, since context-to-context attention in the final layer does not influence the policy output. Across layers, attention patterns shift from localized to distributed, suggesting that the model first stitches transitions locally before constructing global structure.
Overall, we suggest that the model organizes memory to reflect Euclidean structure, using that geometry to guide action selection. This interpretation is supported by the model’s ability to take spatial shortcuts through previously unseen states (Fig. 2C).
Tree Mazes In tree mazes, a useful strategy can be to identify when the agent is on a critical path to reward and to default to the parent-node action otherwise. This is because there are only $6$ states in the maze (of $127$) where the optimal action is to transition to the left child or right child (Fig 7B). These are the states on the path from root to reward, which we will call the left-right (L-R) path. Indeed, the model has a strong action bias to take parent-node transitions (Appendix N). Overall, we find evidence that the model exploits this structure and learns the following strategy:
@@
We arrive at this hypothesis by asking how the model takes correct actions when it is on the L-R path. Again, we work backwards and analyze which context-memory tokens are sufficient to influence the output from the final model layer. We do so by forcing the query token to attend only to select context-memory tokens (Fig 7C). Surprisingly, we find that the model output is unaffected if the query token of the last layer attends only to context-memory tokens involving the query state. These tokens contain sufficient information for the model to make its decision.
@@
With this in mind, our next question was to understand what information is contained in the context-memory tokens entering the last model layer. We repeat our linear decoding analyses on the context-memory tokens. Two variables are well-decoded. First, the inverse action for the transition represented in a context-memory token can be decoded with high accuracy (Fig 7D). The other well-decoded variable is whether the context-memory token is a transition on the L-R path (Fig 7D), regardless of direction (i.e., towards or away from goal). Possibly, at decision time the model tests if there are context-memory tokens that contain the query state and are tagged as being on the L-R path. If so, then the correct left/right action can be inferred from the same tagged tokens (in particular since inverse actions are also encoded). We find further evidence for this strategy by re-doing our sufficiency analysis from Fig 7C. We test three additional restrictions to the previous analysis: (1) tokens must be on the L-R path in the direction to the goal, (2) tokens must be on the L-R path in the direction away from the goal, or (3) tokens are not on the L-R path at all. We find that in the first two cases the model output is unaffected, but in the third case the output is negatively impacted (Appendix N). The results of these perturbations are consistent with our hypothesis that decision-making in the model relies on identifying if the query state is on the L-R path via intermediate computations stored in context-memory tokens.
@@
Finally, we ask how this information becomes present in the context-memory tokens. We plot the L-R path decoding accuracy from Fig 7D by the distance from the context-memory token to reward (Fig 7E). Across model layers, we see that the decoding accuracy first improves for tokens closest to the goal and for tokens farthest from the goal. The accuracy then improves for tokens at an intermediate distance from the goal, where more transition information must be integrated to know if the token is on the L-R path. Analysis of context-to-context attention also shows more integration of local information in the first model layer than in the second (Fig 7F). Taken together, we suggest that path stitching occurs between context-memory tokens such that L-R path tokens are tagged expanding outwards from the reward state.
@@
Figure 7: Tree mazes are solved by tagging context-memory tokens on a critical path to reward. A. Overview of model analysis methods used in tree maze B. We ask how the model correctly takes action on the path from root to goal (L-R path), the only states in the maze where the optimal action is to transition to the left or right child. C. Change in cross-entropy loss when attention is fixed to only the context-memory token of interest, plotted by token position. In all tests, the query state remains the same. Line color indicates layer of intervention. Mean across $50$ environments, shaded regions show $95\%$ confidence interval. D. Balanced accuracy of decoding from context-memory tokens their presence on the L-R path and the inverse action. Decoding inverse action is only non-trivial when considering parent-node transitions so we test our decoder only on those transitions. Mean across $60$ environments with error bars showing $95\%$ confidence interval. We decode from the input context-memory tokens for each model layer (line color indicates layer). Dashed line indicates chance. E. L-R Path decoding accuracy from (D), but separated by how far each context-memory token is from goal. F. As in Fig 6 but for tree mazes.
@@ 5 Conclusion
We have shown that rapid adaptation of agents in tasks relevant to natural cognition can be explained by RL strategies that lie outside traditional model-free or model-based frameworks. Despite this, our meta-learned model also displays phenomena expected from neural activity: learning of environment structure, alignment of representations across environments, and biased memory recall patterns at decision-time. Taken together, this suggests that understanding the cognitive processes that support rapid learning may require theorists to consider a broader space of planning strategies. Finally, our analysis of the RL strategies that emerge in transformers suggest a novel use of episodic memory– each memory is not only a record of the original experience, but also stores additional computation useful for decision making [Dasgupta and Gershman, 2021].
6 Acknowledgements
We thank Ekdeep Lubana, Jack Lindsey, Gautam Reddy, Ben Eysenbach, Lily Zhang, and members of the Rajan lab for helpful discussions and feedback. This work was supported by the Harvard Medical School Lefler Small Grant Award and Dean’s Innovation Award, NIH (RF1DA056403), James S. McDonnell Foundation (220020466), Simons Foundation (Pilot Extension-00003332-02), McKnight Endowment Fund, CIFAR Azrieli Global Scholar Program, and NSF (2046583).
Appendix A Task construction
A.1 Gridworld
@@
We use a $5\times 5$ 2D gridworld environment. Thus, there are $N=5\times 5=25$ states in the environment, each of which corresponds to an underlying $(x,y)$ location. Actions are one-hot encoded and consist of: up, right, down, left, stay. If the agent chooses to take an action that hits the environment boundaries, this manifests as a "stay" transition. The transition structure in this environment $\mathcal{T}$ (that is, how actions transition the agent from one $(x,y)$ state to another) is fixed across all tasks. Each task is defined by the sensory encoding $\mathcal{S}$, the reward location $\mathcal{R}$, and the in-context exploration trajectory $\mathcal{D}$.
Each state in a gridworld task corresponds is encoded by a $10$-dimensional vector. For each task, we describe the set of these $N$ encoding vectors as $\mathcal{S}$. The following describes how we generate the encoding vectors comprising $\mathcal{S}$. We first define a random expansion matrix $E\in\mathcal{R}^{N\times N}$, where $E_{i,j}\in\mathcal{N}(0,1)$. We next construct a distance correlation matrix $D\in\mathcal{R}^{N\times N}$ by exponentiating the negative Euclidean distances between all pairs of grid positions: $D_{i,j}=\sigma^{(||(x_{i},y_{i})-(x_{j},y_{j})||_{2})}$ for states $i$ and $j$ and their corresponding $(x,y)$ locations. Here, $\sigma\in[0,1]$ is a correlation parameter that controls how strongly nearby positions are correlated in the encoding space. Thus, the encoding of state $i$ is computed as $\frac{ED_{:,i}}{||ED_{:,i}||_{2}}$ and $\mathcal{S}=\{\frac{ED_{:,i}}{||ED_{:,i}||_{2}}\}_{i=0}^{N}$.
@@
The reward state $\mathcal{R}$ is chosen from the $N$ states in the environment. The in-context exploration trajectory is generated from a random walk with a randomly chosen initial state, plus some reasonable heuristics. Specifically, we make the probability of selecting the "stay" action half as likely as the other actions. In addition, if the agent takes an action that causes it to not transition to a new state, the probability of taking that action again is downweighted to 00 until the agent transitions to a new state (preventing the agent from getting stuck at boundaries). Running this biased random walk for $T$ steps gives us $C={(s_{t},a,s^{\prime}_{t},r_{t})}_{t=0}^{T}$, a set of standard RL transitions.
@@
A.2 Tree Mazes
We use binary tree environments with $7$ tree layers (that is, a minimum of $6$ actions is needed to navigate from root to leaf). Actions are one-hot encoded and consist of: right child, left child, parent, stay. If the agent tries to transition to a node that does not exist (e.g. trying to go to "parent" from the root), this manifests as a "stay" transition. The transition structure $\mathcal{T}$ can vary across tasks. This is because in each task the underlying tree is generated with branching probability $0.9$. Thus, there is a maximum of $127$ states in each task. Each task is defined by the transition structure $\mathcal{T}$, sensory encoding $\mathcal{S}$, the reward location $\mathcal{R}$, and the in-context exploration dataset $\mathcal{D}$.
The sensory encodings are generated as in gridworld, except $D$ is defined via the geodesic distances between any two tree states. The reward state $\mathcal{R}$ is chosen only from leaf nodes. The in-context exploration trajectory $\mathcal{D}$ is generated from a random walk from the root node, with reasonable heuristics. We use the same heuristics as in gridworld. We also add heuristics described in Rosenberg et al. [2021] of mice in similar mazes. That is, the agent is more likely to alternate between left and right transitions when transitioning through child nodes. In addition, the agent is less likely to backtrack.

### v3

@@ 3 Results
Gridworld A key signature of cross-context structure learning is the alignment of internal representations across environments with shared topology: even when sensory observations differ across environments, states occupying the same grid location should be encoded more similarly than states from different locations (Fig. 4A). To test this, we compute pairwise correlations between model representations of the same graph node across different test environments (Fig. 4B). We separate correlation scores by whether the compared states occupy the same or different graph nodes. With limited context, representations appear collapsed—showing high correlation across all states regardless of node identity. As context length increases, this gap widens: states from the same node become more aligned than those from different nodes (Fig. 4B; solid vs. dashed).
@@
Tree Mazes We repeat the same analyses in tree mazes and find similar alignment strategies emerge (Fig. 4DE). We further analyze the correlation scores by the node position within the tree (Fig. 4F). Cross-context alignment is strongest for states near the root or leaves of the tree. This is consistent with §3.2, where the representations capture coarse structure over precise positional detail in tree mazes. Importantly, neither in-context nor cross-context representation alignment was explicitly trained—these strategies emerge opportunistically as a byproduct of meta-learning.
@@
Figure 5: Memory retrieval at decision time shows limited expansion from the query state and the goal state. A. Example tree maze environment, context length 800. Edge color indicates gradient attribution strength for each transition. B. Average attribution strength of each context memory vs distance from query state and goal state, across 50 environments.
3.4 RL strategy is neither value-based learning nor model-based planning.
Thus far, we have shown that meta-learned agents replicate the rapid learning dynamics observed in animal behavior. A key component of this success is the emergence of structured representations from contextual input. We now turn to characterizing the mechanisms of the underlying RL strategy. We begin by testing whether the model exhibits hallmarks of standard model-free reinforcement learning. We test whether value information can be linearly decoded from model representations, but overall did not find evidence for this (Apps. K and L).
[…]
@@
In summary, we find that the agent does not use value gradients or path planning to make decisions. We suggest that the learned in-context strategy lies outside the standard taxonomy of model-free and model-based reinforcement learning. Our analyses also reveal an additional neural prediction: memory retrieval at decision time should be biased toward experiences near the agent’s current location and its goal. Such replay patterns have been observed in the hippocampus during spatial decision-making tasks (jackson2006hippocampal; pfeiffer2013hippocampal; mattar2018prioritized).
@@
Figure 6: Gridworld tasks are solved by aligning internal representations to Euclidean space. A. Overview of analysis steps. B. Total XY decoding error across $60$ test environments, plotted by true XY position of the query state, given query token embeddings from layer 1 (left) and layer (right). Context length = 250. C. As in (B), but for decoded vs. true angle from query state to the goal. D. Change in cross-entropy loss after ablating context tokens along the query–goal path, plotted by ablated token position. Line color indicates layer of intervention. E. Average attention score between context-memory tokens as a function of spatial distance between tokens, for layers 1 and 2. (D,E) show mean over 50 environments, with shading for 95% C.I.
@@
3.5 Models learn strategy where intermediate computations are stored in context-memory tokens.
@@
    Use in-context experience to align representations to Euclidean space.
@@
We arrived at this hypothesis by first identifying the task-relevant variables that can be linearly decoded from the query token at each layer (Fig. 6A). We train a linear decoder to predict the underlying XY position of the query state from its embedding (App. K) and evaluate accuracy on held-out environments with novel sensory observations. Decoding accuracy improves across layers, with spatial position becoming reliably recoverable by layer 2 (error <1<1; Fig. 6B). Building on this spatial structure, we next decode the angle from the query state to the goal– again, this information can be accurately decoded by layer 2 (Fig. 6C). Interestingly, both XY position and angle-to-goal can be decoded from the embeddings of the context-memory transitions $(s,a,s^{\prime},r)$ as well.
To localize where angle-to-goal information may be computed, we test which context-memory tokens are necessary for correct decisions. Using attention ablations, we show that model performance relies on attending to tokens near the query and goal states in layer 2 (Fig. 6D, App. N). We propose that layer 2 extracts the internal XY coordinates from query and goal state tokens to compute the relative angle between them. To understand how XY information arises, we examine how state representations evolve through context-to-context attention.
We show that, across layers, attention patterns between context memory tokens shift from localized to distributed (Fig. 6E, App. N), suggesting that the model first stitches transitions locally before constructing global structure. Overall, we suggest that the model organizes memory to reflect Euclidean structure and use that geometry to guide action selection. This explains the model’s ability to take unseen shortcuts (Fig. 2C).
Tree Mazes In tree mazes, a useful strategy can be to identify when the agent is on a critical path to reward and to default to the parent-node action otherwise. This is because there are only $6$ states in the maze (of $127$) where the optimal action is to transition to the left child or right child (Fig 7B). These are the states on the path from root to reward, which we will call the left-right (L-R) path. Indeed, the model has a strong action bias to take parent-node transitions (App. O). Overall, we find evidence that the model exploits this structure and learns the following strategy:
@@
We arrive at this hypothesis by asking how the model takes correct actions when it is on the L-R path. Again, we work backwards and analyze which context-memory tokens are sufficient to influence the output from the final model layer (Fig 7C, App. O). Surprisingly, we find that the model output is unaffected if the query token of the last layer attends only to context-memory tokens involving the query state. These tokens contain sufficient information for the model to make its decision.
@@
With this in mind, our next question was to understand what information is contained in the context-memory tokens entering the last model layer. We repeat our linear decoding analyses on the context-memory tokens. Two variables are well-decoded. First, the inverse action for the transition represented in a context-memory token can be decoded with high accuracy (Fig 7D). The other well-decoded variable is whether the context-memory token is a transition on the L-R path (Fig 7D), regardless of direction (i.e., towards or away from goal). Possibly, at decision time the model tests if there are context-memory tokens that contain the query state and are tagged as being on the L-R path. If so, then the correct left/right action can be inferred from the same tagged tokens (in particular since inverse actions are also encoded). We find further evidence for this strategy by re-doing our sufficiency analysis from Fig 7C with additional restrictions on the selected tokens (see App. O).
@@
Finally, we ask how this information becomes present in the context-memory tokens. We plot the L-R path decoding accuracy from Fig 7D by the distance from the context-memory token to goal (Fig 7E). Across model layers, we see that the accuracy first improves for tokens closest to or farthest from the goal. Later, the accuracy improves for tokens at an intermediate distance from the goal, where more transition information must be integrated to know if the token is on the L-R path. Furthermore, attention patterns between context memory tokens shift from localized to distributed across model layers (Fig 7F). Taken together, we suggest that path stitching occurs between context-memory tokens such that L-R path tokens are tagged expanding backwards from the goal.
@@
Figure 7: Tree mazes are solved by tagging context-memory tokens on a critical path to reward. A. Overview of analysis steps. B. We focus on the path from root to goal (L-R path). C. Change in cross-entropy loss when attention is fixed to context-memory tokens at specific points on the L-R path. In all tests, the query state is the root. Line color indicates layer of intervention. Mean across $50$ environments, shading shows $95\%$ C.I. D. Given embeddings of context-memory transitions, the balanced accuracy of decoding their presence on the L-R path and the inverse action. Decoding the inverse action is only non-trivial for parent-node transitions so we test only on those. Embeddings are taken from the input tokens to each model layer (line color). Mean across $60$ environments, error bars show $95\%$ C.I. Dashed line is chance. E. L-R Path decoding accuracy from (D), but separated by how far each context-memory transition is from goal. F. As in Fig 6E but for tree mazes.
@@ 4 Conclusion
We have shown that rapid adaptation of agents in tasks relevant to natural cognition can be explained by RL strategies that lie outside traditional model-free or model-based frameworks. Despite this, our meta-learned model also displays phenomena expected from neural activity: learning of environment structure, alignment of representations across environments, and biased memory recall patterns at decision-time. Taken together, this suggests that understanding the cognitive processes that support rapid learning may require theorists to consider a broader space of planning strategies.
Finally, our analysis of the RL strategies that emerge in transformers suggest a novel use of episodic memory– each memory is not only a record of the original experience, but also stores additional computation useful for decision making (dasgupta2021memory).
Reproducibility
Code will be publicly available in a Github link in the final paper (after de-anonymization).
Appendix A Related Works
Meta-learning to discover in-context reinforcement learning algorithms In order to develop a model that can in-context reinforcement learn, we use a recently introduced meta-learning framework (lee2023supervised). Meta-learning is concerned with “learning-to-learn”, using prior knowledge learned from previous tasks to support rapid adaptation to new ones (beck2023survey). That is, the goal is to learn an algorithm $f$ that can be deployed in new tasks. The learning of $f$ is called the outer-loop while $f$ itself is referred to as the inner-loop (beck2023survey). In many settings where $f$ is a RL algorithm (meta-RL), the outer loop shapes the weights of the network, but learning within each task occurs via activation dynamics– through memory and internal state– rather than parameter updates (beck2023survey; lampinen2024broader). The resultant $f$ is considered an in-context learning algorithm (beck2023survey; sandbrink2024modelling; lampinen2024broader).
Early examples of this approach include RL2, which meta-trains a recurrent neural network using reinforcement learning in the outer loop, such that an in-context RL strategy emerges in the inner loop (wang2016learning; duan2016rl). Subsequent work extended this approach to include explicit episodic memory mechanisms, combining RNNs with key–value memory architectures (ritter2018been; ritter2020rapid; team2023human). More recently, lee2023supervised proposed decision-pretrained transformers (DPTs), which use supervised training in the outer loop to induce in-context RL behavior in the inner loop. We adopt DPTs in this work both for practical reasons–their scalability and ease of training–and for scientific ones: the transformer architecture allows us to probe how memory-based computation supports in-context learning.
Meta-learning to describe cognition and neural activity Meta-learning has been proposed as a framework to model both cognitive flexibility and structured learning in neuroscience and psychology (binz2024meta). In human cognitive tasks, meta-learned models have been used to replicate observed heuristics in decision-making and to account for few-shot generalization (dasgupta2020theory; binz2022heuristics; lake2023human). Meta-RL, in particular, has been used to generate hypotheses about how neural systems implement learning across tasks. Previous studies have used the RL2 framework to show how the outer and inner learning loops can model different areas of the brain, with the prefrontal cortex often playing a key role (wang2018prefrontal; hattori2023meta; zheng2025flexible). Despite these advances, many computational models in neuroscience rely on single-task RL, training agents independently on each task without leveraging prior experience. In contrast, we use meta-RL as a tool for developing flexible in-context learning, without attempting to localize the outer loop to any specific brain region. Our focus is on the computational content of the learned representations and decision-making strategies.
Transformers and episodic memory systems Transformer models process sequences by computing self-attention over key–value pairs, enabling flexible access to information across long contexts (vaswani2017attention). This key–value structure has led to interpretations of transformers as memory systems (geva2020transformer), aligning them with a broader class of models that incorporate explicit memory mechanisms (graves2014neural; sukhbaatar2015end; graves2016hybrid; banino2020memo). These systems separate memory addressing (keys) from memory content (values), enabling high-fidelity storage and targeted retrieval. This architectural separation bears many similarities to theoretical accounts of episodic memory in the brain (teyler1986hippocampal; teyler2007hippocampal; gershman2025key). For example, recent work has formalized connections between key–value architectures and Hopfield networks, a classic model of associative memory in the brain (krotov2020large; ramsauer2020hopfield). Related approaches such as fast-weight models (ba2016using; munkhdalai2017meta) offer alternative mechanisms for temporary memory storage and in-context computation, often drawing from Hebbian or synaptic dynamics. Other studies have proposed biologically grounded implementations of key–value attention mechanisms, further linking transformer-like architectures to neural computation (bricken2021attention; tyulmankov2021biological; whittington2021relating; kozachkov2023building; fang_2025; chandra2025episodic). Several of these models take direct inspiration from the hippocampus, a brain region widely implicated in episodic memory (whittington2021relating; fang_2025; chandra2025episodic). Recent experimental work has also identified key–value–like coding patterns in hippocampal activity during episodic memory tasks (chettih2024barcoding). Together, these findings suggest that key–value architectures offer a useful computational analogy for episodic memory. In this work, we analyze a meta-trained transformer to examine what kinds of memory-supported strategies emerge when such an architecture is optimized for rapid adaptation.
Appendix B Task construction
B.1 Gridworld
@@
We use a $5\times 5$ 2D gridworld environment. Thus, there are $N=5\times 5=25$ states in the environment, each of which corresponds to an underlying $(x,y)$ location. Actions are one-hot encoded and consist of: up, right, down, left, stay. If the agent chooses to take an action that hits the environment boundaries, this manifests as a ”stay” transition. The transition structure in this environment $\mathcal{T}$ (that is, how actions transition the agent from one $(x,y)$ state to another) is fixed across all tasks. Each task is defined by the sensory encoding $\mathcal{S}$, the reward location $\mathcal{R}$, and the in-context exploration trajectory $\mathcal{D}$.
[…]
@@
The reward state $\mathcal{R}$ is chosen from the $N$ states in the environment. The in-context exploration trajectory is generated from a random walk with a randomly chosen initial state, plus some reasonable heuristics. Specifically, we make the probability of selecting the ”stay” action half as likely as the other actions. In addition, if the agent takes an action that causes it to not transition to a new state, the probability of taking that action again is downweighted to 0 until the agent transitions to a new state (preventing the agent from getting stuck at boundaries). Running this biased random walk for $T$ steps gives us $C={(s_{t},a,s^{\prime}_{t},r_{t})}_{t=0}^{T}$, a set of standard RL transitions.
@@
B.2 Tree Mazes
We use binary tree environments with $7$ tree layers (that is, a minimum of $6$ actions is needed to navigate from root to leaf). Actions are one-hot encoded and consist of: right child, left child, parent, stay. If the agent tries to transition to a node that does not exist (e.g. trying to go to ”parent” from the root), this manifests as a ”stay” transition. The transition structure $\mathcal{T}$ can vary across tasks. This is because in each task the underlying tree is generated with branching probability $0.9$. Thus, there is a maximum of $127$ states in each task. Each task is defined by the transition structure $\mathcal{T}$, sensory encoding $\mathcal{S}$, the reward location $\mathcal{R}$, and the in-context exploration dataset $\mathcal{D}$.
The sensory encodings are generated as in gridworld, except $D$ is defined via the geodesic distances between any two tree states. The reward state $\mathcal{R}$ is chosen only from leaf nodes. The in-context exploration trajectory $\mathcal{D}$ is generated from a random walk from the root node, with reasonable heuristics. We use the same heuristics as in gridworld. We also add heuristics described in rosenberg2021mice of mice in similar mazes. That is, the agent is more likely to alternate between left and right transitions when transitioning through child nodes. In addition, the agent is less likely to backtrack.